# INRetouch: Context Aware Implicit Neural Representation for Photography Retouching

### v3

@@ 5 Experiments
For the training set, we used 510 images with 150 different presets. For the test set, we used a new unseen 22 presets with the remaining 60 images. For our proposed method, we learn directly from the reference, so we only used the testing setup.
@@
As mentioned in the dataset section 3, the same preset can generate a different-looking output when applied to different images. To make sure that the difference between the input and the GT visually matches the difference in the reference pair, for each input image we choose the reference that closely matches the color distribution of the input. We do so by comparing the 3D color histogram between the input image and the reference before edit. For a fair comparison, we use the same reference for all the tested models. We use PSNR and SSIM to evaluate the accuracy of the retouch transfer.
@@
We train our INR architecture for 1000 iterations with a sampling window size of 13 and 484 samples per iteration, using the L1 loss function. We use Adam [29] optimizer with a learning rate of $1e^{-3}$ that is gradually decreased to $1e^{-4}$ using Cosine Annealing [37] learning rate scheduler. We provide more technical details related to data pre-processing, training, and reproducibility in the supplementary material.
@@ 5.1 Retouching Transfer Benchmark
Table 1: Performance of different models in Retouch Transfer. Our method performs the best in retouch transfer with learning only from the reference sample. Our method is by far the simplest and most efficient while achieving the best results.
Type Method PSNR $\uparrow$ SSIM $\uparrow$
Full Data Training StyleGan [26] 20.6370 0.7587
Deep Preset [23] 21.9494 0.7727
Neural Preset [27] 22.1291 0.7602
Example Based (No Training) Image Analogies [22] 12.3230 0.4031
Deep Image Analogies [35] 12.7634 0.3195
Painter [55] 12.2027 0.3500
Visual Prompting [5] 14.6115 0.4129
INR (One Shot) LTE [7] 16.2378 0.6090
CiaoSR [7] 19.1142 0.6936
LIT [10] 18.5052 0.6558
InRetouch(Ours) 23.4216 0.8054
Figure 4: Comparison between different methods on retouching transfer task. Our method learns the edits effectively from a single sample generalizing to wide variety of edits and has the most consistent output with the GT. We can appreciate the ability of our method to learn and adapt to complex edits like venting and local modification.
@@
Deep Preset[23] 36.6 512.14 0.0578 1160.84 0.1542 2048.53 0.2820 4609.19 0.6601 21.9494
StyleGan [26] .61524 24.86 0.0316 55.95 0.0681 99.4 0.1380 223.78 0.2926 20.6370
Neural Preset [27] 4.8 153.87 0.0902 348.77 0.2355 615.47 0.4293 OOM OOM 22.1291
SIREN [48] 0.00800 8.2 0.0055 18.45 0.0124 32.81 0.0191 73.81 0.0459 22.8655
SIREN-Split [32] 0.01085 10.0 0.0087 22.5 0.0219 40.0 0.0301 90.0 0.07617 23.1025
@@
For a comprehensive benchmark on retouch transfer, we tested different neural network architectures that were proposed for generative and style transfer tasks, to show their performance on the newly proposed task. For an accurate comparison, we adapted these networks to work with our proposed task and were trained on our dataset. To ensure accurate training, we chose references that match the edits between the input and GT, similar to the evaluation process 5. Additionally, we included other example-based methods that require no training and were developed for general-purpose example-based applications. We tested 2 kinds of these approaches, including image analogies [22, 35] and In-Context learning [55, 5] approaches. Lastly, we tested different INR architectures that were proposed for image restoration tasks. For a fair comparison, we only used the INR architecture without the encoder and we used the same pipeline as our method. It is important to highlight that all the methods compared have the same input information.
@@
As we see in Tab. 1 we achieve the best performance with a big margin over the other methods. Methods that require training on the full dataset produce big discrepancies in performance between seen and unseen styles as they struggle to generalize for new unseen styles.
The methods that require no training struggle in the retouch transfer task. The image analogies method works by copying from the reference, and is limited to the information in the reference. The In-Context learning methods fail to recognize the required task from the given reference.
The tested INR architectures include complex parameter intensive modules (self-attention[16]). Because of the complex architecture, they fail to generalize to new input because of the limited training samples (single reference).
Our proposed method performs the best, learning only from the given reference which allows it to generalize to edits and styles without depending on the available data variety. Additionally, our method is by far the most efficient with only 11.5 K parameters which makes it very practical and allows for high-resolution image editing in very limited hardware. Our extensive experiments show that the simplicity of our proposed architecture without any complex or parameter-intensive layers is crucial to learn from a single sample and generalize to new inputs.
Figure 5: Importance of context awareness for local and region specific modifications. We can appreciate our method ability to recognize different objects and apply local and region-specific edits accurately.
@@
In Fig. 4 we show the quality of our output compared to other methods. We can appreciate the consistency of our method with the ground truth producing a high-quality output without artifacts. We can notice the ability of our method to learn complex edits accurately like the smooth vignetting effect (row 1), and can produce local and content-based edits accurately (Sky in row 2). Methods that require full training are limited to the training dataset and might fail to generalize to new edits (row 2,3). Additionally, we can notice generated artifacts in their outputs failing to apply smooth edits and high-quality output.
In-context learning methods such as Painter [55], fail to recognize the required task, and Image Analogies [35] are limited to the reference information, producing unreliable output. The Other INR method CiaoSR [7], tends to overfit on the reference, failing to generalize to new input. Our proposed method is able to overcome the limitations of previous methods.
@@
As we see in Tab. 2, our proposed method is very efficient with very few parameters (11k) requiring very little memory and with a fast inference time, processing a 4K image in just 70 ms. The style transfer methods compared (first 3 rows), have much more parameters (60x to 400x more parameters) and need 5x to 10x more time to process the same image. Additionally, in comparison to the INR methods we can notice our model achieving a similar efficiency and inference speed. Even though our new INR method requires more parameters, our model design and layer choices were able to maintain a similar efficiency while noticeably improving the performance.
@@
Including neighboring pixels allows the model to recognize texture, edges, and context, which is important information to apply region-specific and object-specific modification. As we see in Fig. 5 ordinary pixel-wise INR architecture fails to recognize objects and regions and fails to transfer local modification to the new input. Our method is able to recognize the objects (row 1, 3) and apply separate edits to them. In the first example, our method is able to place fog around the center object accurately, while the ordinary MLP overfits on the reference. Additionally, our method is able to simulate operations like blurring (row 2) because of the access to neighboring pixels.
@@
Designed to be lightweight and efficient, our model enables affordable inference, which motivated us to extend its application to video editing. As shown in the accompanying video in supplementary materials, our method effectively learns edits from images and applies them to videos, producing visually pleasing results with excellent temporal consistency and no noticeable artifacts. This can be attributed to the editing clarity from the use of before and after editing reference and designing our method to focus on color modification through local awareness.
Unlike existing methods, such as style transfer and generative-based models, which often struggle with temporal consistency and introduce significant noise, our approach overcomes these limitations. This demonstrates both the effectiveness of our network and the controllability of the learned edits.
@@ 5.2 Context-aware INR for Image Reconstruction
To show the importance of context awareness, we test it on a variety of INR image reconstruction applications. We tested on Gamut Mapping [31], Metadata-Based RAW Reconstruction [32], Neural Implicit LUT [14].
@@
Gamut Mapping[31] SIREN 54.4262 0.9997
@@
Neural LUT[14] SIREN 33.9761 0.9632
@@
In Tab. 3 We show the difference between traditional INR architecture with pixel sampling and MLP layers, in comparison with our method that employs window sampling and context awareness. The difference between the 2 tested architectures is the sampling technique and the addition of context awareness. As we see in Tab. 3, our proposed INR architecture consistently performs better, showing the importance of context for image-related tasks. Our method proves effective in different tasks while maintaining the advantages of INR for efficiency and speed.
5.3 Context-Aware INR Ablation
@@
Figure 6: The effect of different components in our architecture on the output.
@@
Convolution [30] 23.4348 0.8091 47.78 44.03
Deform convolution [64] 20.9221 0.7470 24.36 63.29
Self-Attention [16] 23.1863 0.8000 19.17 18.74
Depth wise Convolution [13] 23.4216 0.8054 11.49 10.59
INR Components Ablation
In Tab. 4 we show the performance improvement of changes we made over the ordinary MLP INR architecture. For our proposed architecture, processing color and coordinates separately proved effective when giving more attention to the pixel color by using different regularization weights. Context awareness through the Depth-Wise Conv layer achieves the biggest improvement proving the importance of neighboring information in our task. As we see in Fig. 6, context awareness enables recognizing textures, objects, and edges, producing less artifacts and better editing.
@@
In Tab. 5 We tested different layers to process neighboring pixels in our INR architecture. Complicated and parameter intensive layers like deform convolution and self-attention perform worse because of the limited training samples (one sample) which doesn’t allow to generalize to new input images. For our final model, we chose the depth-wise convolution layer as it is the most efficient without performance loss, maintaining the speed and efficiency of traditional INR.
@@
w/ Fourier Features [52] 21.5339 0.7537 11.49 10.59
w/ RDN Features [62] 23.1843 0.7961 22140 20278
w/ SWINIR Features [34] 22.9198 0.7875 11770 10685
@@
A common practice is to encode the input information to the INR to generate a more expressive input. We tested different kinds of encoding using Fourier encoding and pre-trained feature extractor. We tested a CNN [62] and Transformer [34] based features extractors pre-train on 2X image super resolution task. As we see in Tab. 6 using the RGB values directly without encoding archives the best performance. In our task, INR tends to overfit on the reference when using input encoding. Although feature extractors increase the receptive field over the input information and help the INR process the input information, it requires the INR to decode these features. Additionally, it adds a huge computational cost to the pipeline.
Number of References
As we see in Fig. 7 (left) increasing the number of references improves the performance, which shows the ability of our method to interpolate and extract information from multiple sources. This is effective when working with predefined styles with multiple references available.
Convergence Speed
Our window sampling technique proves effective as we see in Fig. 7 (right), achieving a similar optimization speed as pixel sampling because of the flexibility with the number of updates. Additionally, it allows the integration of context awareness, improving the performance. We can also notice the issue of full image training as it requires more time to optimize.
@@
As mentioned in Section 3, one of the biggest challenges with the dataset is the inconsistency with the preset output. This inconsistency is also an issue of presets in photo editing software for the end user. They provide inconsistent visual edits when using the same preset across different images, which limits their editing transfer capabilities. Our method can serve as an alternative to presets for edit transfer providing more consistent visual edits – see Fig. 8. It can generalize to new edits as it learns adaptively from the given reference. Additionally, it will provide a platform-independent method for edits transferring between different editing software.
@@ 6 Conclusion
We propose a novel image enhancement problem focused on image editing and retouching images. First, we present a novel dataset for image retouching that poses new challenges and improves previous datasets such as MIT-5K. Second, we propose a novel method that employs context-aware implicit neural representations (INRs) for learning complex image editions. Our results show the potential of our approach as a general image retouching method, improving the performance of INR on different image restoration tasks. Moreover, generative methods for image editing could also benefit from our dataset and task. Our code and datasets will be open-source upon acceptance.
Looking forward, this work opens new possibilities for research in adaptive image manipulation, suggesting that complex editing operations can be effectively learned and transferred without sacrificing quality or control. We believe our contribution provides a foundation for future work in automated photo editing.
@@
We first kindly refer the readers to our accompanying video examples.
Then, in this supplementary material, we provide more implementation details of our work in Sec. 1. We also provide more ablation studies in Sec. 2.
As for visuals, we first provide a comparison on visual consistency in Sec. 3. More visuals on retouching transfer comparison can be found in Sec. 4. Finally, in Sec. 5, we show the variety of our presets applied to a natural image.

### v4

@@ 5 Experiments
For the training set, we used 510 images with 150 different presets to train the compared methods. For the test set, we used a new unseen 22 presets with the remaining 60 images. For our proposed method, we learn directly from the reference, so we only used the testing setup.
@@
As mentioned in the dataset section 3, the same preset can generate different-looking outputs when applied to different images, because many edits depend on color distribution (e.g. highlight, shadow, hue, etc) and local properties. Therefore, the optimal evaluation setup involves using a reference (before-after) where the input (before) image has similar properties as the testing input image. To achieve this, we compare the 3D color histogram between the input image and the reference before editing. This ensures that for each particular testing image, the approaches are tested in an optimal way. For a fair comparison, we use the same reference for all the tested models.
@@
We train our INR architecture for 1000 iterations with a sampling window size of 13 and 484 samples per iteration, using the L1 loss function. We use Adam [kingma2014adam] optimizer with a learning rate of $1e^{-3}$ that is gradually decreased to $1e^{-4}$ using Cosine Annealing [loshchilov2016sgdr] learning rate scheduler. We use a single context awareness block for our final architecture. We provide more technical details in the supplementary
@@ 5.1 Retouching Transfer Benchmark
Table 1: Performance of different models in Retouch Transfer. Our method performs the best with learning only from the reference sample. Our method is by far the simplest and most efficient while achieving the best results.
Type Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPs $\downarrow$
Full Data
Training
StyleGan [karras2019style] 20.6370 0.7587 0.1958
Deep Preset [ho2021deep] 21.9494 0.7727 0.1868
Neural Preset [ke2023neural] 22.1611 0.7690 0.1763
Example Based
(No Training)
Image Analogies [hertzmann2023image] 12.3230 0.4031 0.7693
Deep Image Analogies [liao2017visual] 12.7634 0.3195 0.7472
Painter [wang2023images] 12.2027 0.3500 0.7798
Visual Prompting [bar2022visual] 14.6115 0.4129 0.6326
INR
(One Shot)
LTE [yang2020learning] 16.2378 0.6090 0.3763
CiaoSR [cao2023ciaosr] 19.1142 0.6936 0.2275
LIT [chen2023cascaded] 18.5052 0.6558 0.3099
InRetouch(Ours) 23.4216 0.8054 0.1490
@@
Deep Preset[ho2021deep] 36.6 512.14 0.0578 1160.84 0.1542 2048.53 0.2820 4609.19 0.6601 21.9494
StyleGan [karras2019style] .61524 24.86 0.0316 55.95 0.0681 99.4 0.1380 223.78 0.2926 20.6370
Neural Preset [ke2023neural] 4.8 153.87 0.0902 348.77 0.2355 615.47 0.4293 OOM OOM 22.1291
SIREN [sitzmann2020siren] 0.00800 8.2 0.0055 18.45 0.0124 32.81 0.0191 73.81 0.0459 22.8655
SIREN-Split [li2023metadata] 0.01085 10.0 0.0087 22.5 0.0219 40.0 0.0301 90.0 0.07617 23.1025
@@
Figure 4: Comparison between different methods on retouching transfer task. Our method learns the edits effectively from a single sample, generalizing to a wide variety of edits, and has the most consistent output with the GT. We can appreciate the ability of our method to learn and adapt to complex edits like vignetting and local modification.
For a comprehensive benchmark on retouch transfer, we tested different neural network architectures that were proposed for generative and style transfer tasks, to show their performance on the newly proposed task. For an accurate comparison, we adapted these networks to work with our proposed task and were trained on our dataset. To ensure accurate training, we chose references that match the edits between the input and GT, similar to the evaluation process 5. Additionally, we included other example-based methods that require no training and were developed for general-purpose example-based applications. We tested 2 kinds of these approaches, including image analogies [hertzmann2023image, liao2017visual] and In-Context learning [wang2023images, bar2022visual] approaches. Lastly, we tested different INR architectures that were proposed for image restoration tasks. For a fair comparison, we only used the INR architecture without the encoder, and we used the same pipeline as our method. It is important to highlight that all the methods compared have the same input information.
@@
As we see in Tab. 1, Methods that require full training tend to overfit and struggle to generalize to new, unseen styles. Meanwhile, example-based methods that require “no training” are sub-optimal. The tested INR architectures include complex parameter-intensive modules (self-attention[dosovitskiy2020image]). Because of the complex architecture, they fail to generalize to new input because of the limited training samples (single reference).
Our proposed method generalizes to edits and styles without depending on the available data variety. Additionally, our method is the most efficient with only 11.5 K parameters, which makes it very practical and allows high-resolution image editing on very limited hardware.
@@
In Fig. 4 we show the quality of our output compared to other methods. We can appreciate the consistency of our method with the ground truth, producing a high-quality output without artifacts. We can notice the ability of our method to learn complex edits accurately (smooth vignetting effect, row 1) and can produce local and content-based edits accurately (Sky in row 2).
Methods that require full training are limited to the training dataset and might fail to generalize to new edits (row 2,3). Additionally, we can notice artifacts in their outputs, failing to apply smooth edits and high-quality output. In-context learning methods such as Painter [wang2023images] fail to recognize the required task. Image Analogies [liao2017visual] are limited to the reference information, producing unreliable output. Other INR method CiaoSR [cao2023ciaosr] tends to overfit on the reference, failing to generalize to new input.
@@
As we see in Tab. 5.1, our proposed method is very efficient with very few parameters (11k) requiring very little memory and with a fast inference time needing only 1.9 s for training, and processing a 4K image in just 70 ms. The style transfer methods compared (first 3 rows) have many more parameters (60x to 400x more parameters) and need 5x to 10x more time to process the same image. Even though our new INR method requires more parameters in comparison to traditional INR methods, our model design and layer choices were able to maintain a similar efficiency while noticeably improving the performance.
Figure 5: Importance of Context Awareness. The results reflect the importance of context awareness for local and region-specific modifications. Our method InRetouch improves significantly on previous INR methods.
@@
Including neighboring pixels allows the model to recognize texture, edges, and context, which is important information to apply region-specific and object-specific modifications. As we see in Fig. 5, ordinary pixel-wise INR architecture fails to recognize objects and regions and fails to transfer local modification to the new input. Our method is able to recognize the objects (row 1, 3) and apply separate edits to them. In the first example, our method is able to place fog around the center object accurately, while the ordinary MLP overfits on the reference. Additionally, our method is able to simulate operations like blurring (row 2) because of the access to neighboring pixels.
@@
Designed to be lightweight and efficient, our model enables affordable inference, which motivated us to extend its application to video editing. Our method effectively learns edits from images and applies them to videos, producing visually pleasing results with excellent temporal consistency and no noticeable artifacts. This can be attributed to the editing clarity by using before and after editing references, and the design that focuses on color modification through local awareness. Unlike existing methods, such as style transfer and generative-based models, which often struggle with temporal consistency and introduce significant noise, our approach overcomes these limitations.
Figure 6: Qualitative Comparisons. We show the ability of our method to transfer retouches from professionally edited images. We can appreciate the visual consistency between the reference and our method output. Our method is able to recognize objects and apply local and region-specific edits accurately.
Testing on Professionally Edited Images
For a more accurate use case, we tested our method on images edited by professionals (i.e. not created using presets) collected from the internet [^2] https://lightroom.adobe.com/learn/discover. We collected 100 pairs with accompanying presets to compare to Lightroom editing transfer. As we see in Fig. 6, our method can transfer a wide variety of edits, including object/region-specific edits, global edits, vignetting, relighting, etc. Additionally, our method produces more visually similar output to the reference as it learn to transfer editing style, not re-applying editing operations as in Lightroom presets. Moreover, since our method directly learn from the given reference without pretraining, it is not limited to the capabilities of Lightroom presets or editing variations in the dataset. We provide more visuals in the supplementary materials.
@@ 5.2 Context-aware INR for Image Reconstruction
To show the importance of context awareness in INRs, we study our approach on a variety of image reconstruction applications: Gamut Mapping [le2023gamutmlp], Metadata-Based RAW Reconstruction [li2023metadata], Neural Implicit LUT [conde2024nilut].
@@
Gamut Mapping[le2023gamutmlp] SIREN 54.4262 0.9997
@@
Neural LUT[conde2024nilut] SIREN 33.9761 0.9632
@@
In Tab. 3 we show the difference between traditional INR architecture with pixel sampling and MLP layers, in comparison with our method. The difference between the tested architectures is the sampling technique and the addition of context awareness. As we see in Tab. 3, our proposed INR architecture consistently performs better, showing the importance of context for image-related tasks. Our method proves effective in different tasks while maintaining the advantages of INR for efficiency and speed.
5.3 Context-Awareness Ablation Study
@@
INR Components Ablation
In Tab. 4 we show the incremental performance improvement over the baseline MLP INR architecture. For our proposed architecture, processing color and coordinates separately proved effective i.e., we give more attention to the pixel color by using different regularization weights. Context awareness provides the biggest improvement, proving the importance of neighboring information in this task.
@@
Convolution [krizhevsky2012imagenet] 23.4348 0.8091 47.78 44.03
Deform convolution [zhu2019deformable] 20.9221 0.7470 24.36 63.29
Self-Attention [dosovitskiy2020image] 23.1863 0.8000 19.17 18.74
Depth wise Convolution [chollet2017xception] 23.4216 0.8054 11.49 10.59
@@
In Tab. 5 we study different types of layers to bring locality into the INR. The deformable convolution and self-attention are sub-optimal under one/few-shot training. For our final model, we chose the depth-wise convolution layer as it is the most efficient without performance loss, maintaining the speed and efficiency.
@@
w/ Fourier Features [tancik2020fourier] 21.5339 0.7537 11.49 10.59
w/ RDN Features [zhang2018residual] 23.1843 0.7961 22140 20278
w/ SWINIR Features [liang2021swinir] 22.9198 0.7875 11770 10685
@@
A common practice is to encode the input information to the INR to generate a more expressive input. We tested different kinds of encoding using Fourier encoding and pre-trained feature extractors. We tested a CNN [zhang2018residual] and a Transformer [liang2021swinir] based feature extractors pre-trained on a 2X image super resolution task. As we see in Tab. 6, using the RGB values directly without encoding achieves the best performance. In our task, INR tends to overfit on the reference when using input encoding. Although feature extractors increase the receptive field over the input, the INR tends to overfit because of the limited data. Additionally, it adds a huge computational cost.
@@
Number of References
As we see in Fig. 7 (left), increasing the number of references improves performance, showing the ability of our method to interpolate and extract information from multiple sources. This is effective when working with predefined styles with multiple references.
Convergence Speed
Our window sampling technique proves effective, as we see in Fig. 7 (right), achieving a similar optimization speed as pixel sampling because of the flexibility with the number of updates. Additionally, it allows the integration of context awareness, improving the performance. We can also notice the issue of full image training, as it requires more time to optimize.
@@
As mentioned in Section 3, one of the biggest challenges with the dataset is the inconsistency with the preset output. This inconsistency is also an issue of presets in photo editing software. They provide inconsistent visual edits when using the same preset across different images, which limits their transfer capabilities. Our method can serve as alternative to presets for edit transfer, providing more consistent visual edits –see Fig. 8– and generalize to new edits as it learns adaptively from the reference. Additionally, our method is platform-independent, transferring edits between different editing software.
@@ 6 Conclusion
We propose a novel image modification problem focused on image editing and retouching. First, we present the Retouch Transfer Dataset, a novel dataset for image retouching that poses new challenges and improves previous datasets. Second, we propose a novel method that employs context-aware implicit neural representation for learning complex image editions from a single reference. The results show the potential of our approach as a general image retouching method, improving the control and fidelity of style-transfer methods and INRs on different tasks. This work opens new possibilities for research in adaptive image manipulation, suggesting that complex editing operations can be effectively learned and transferred without sacrificing quality or control. We believe our contribution provides a foundation for future work in automated photo editing with a dataset that can be beneficial for other image editing methods.
Acknowledgments: This work was supported by The Alexander von Humboldt Foundation.
@@
We first kindly refer the readers to the project website omaralezaby.github.io/inretouch/ for video results, source code, and dataset.
In this supplementary material, we provide more implementation details of our work in 7. We also provide more ablation studies in 8.
As for visuals, we first provide a comparison on visual consistency in 10. In Section 11 we discuss the limitations of the proposed approach. More visuals on retouching transfer comparison can be found in Section 12. Finally, in Section 13, we show the variety of our presets applied to a natural image.